# Visual Privacy Management with Generative AI for Blind and Low Vision People

### v1

@@ 4.3. Practices of GenAI for Sensitive Content
Medical Content. Participants expressed caution when dealing with medical content while mentioned “privacy” as a reason for preferring GenAI tools over traditional practices. P16 mentioned “I avoid GenAI tools for sensitive documents like medical records. I still go for trusted humans (e.g., family or Aira agents under NDA agreements).” In contrast, P17 mentioned using Generative AI for interpreting medical records, such as mamogram and MRI for better explainability and privacy I would call my sister and tell her to ask my brother-in-law, or another appointment. I used Be my AI and ChatGPT to make sense of my report. Honestly, I liked the idea of being private.” Some participants used GenAI tools to read sensitive physical items. P1 shared: “I used AI for my pregnancy test. It said positive several times, and I was stressed. I took more pictures, and then it said negative. I was relieved.” This highlights how participants preferred using AI independently, even when results varied, to maintain emotional and informational privacy in personal situations.
@@
Financial Document. Participants reported adjusting their methods based on perceived sensitivity and situational privacy needs P9 said “For my credit card, I’d rather use a bonded volunteer. but I have used Be My Eyes AI to read a handwritten check in a pinch, but I knew there was a risk—it could have exposed bank info. Sometimes I just need to know what is in the memo line [..] I don’t want to bug anyone. I try to use secure options like Aira’s free call, which is limited.” P19 on the other hand, used Seeing AI to debit card-“One time I used Seeing AI. I needed to make sure I was throwing away the old debit or new. I used it to read the card numbers and confirm which was which.” Some noted that their choice depends on the level of trust and availability.
@@
Digital & Physcial Media Media. Participants also used GenAI tools to navigate digital and potentially sensitive physical content. P11 expressed caution with document uploads, even on trusted platforms: “ am careful with uploading document/photos. particularly Health Equity one. Even with Aira, who has confidentiality agreements, I am picky. I’d rather call my mom or dad. I sometimes use Seeing AI as a last resort.” P20 reflected on using Microsoft Seeing AI to sort mail “Seeing AI helps me sort mail. There might be something very private in that letter, and I wouldn’t even know when I opened it with AI.” This reflects a privacy pragmatism mindset where participants recognize the potential sensitivity of what they are sharing but justify the decision based on context.
4.4. Sociopolitical Decision Making Factors in Adopting GenAI Tool
@@
We observed some decision-making factors of adopting GenAI, which is shaped by broader concerns about access, access, policy assumptions, and representational harms.
@@
Affordability of GenAI. Participants raised concerns about the cost of GenAI technologies and the assumptions behind their pricing. P11 echoed these concerns, criticizing the pricing and accessibility of tools like Envision AI, and noting disappointment in how products are marketed and supported. P11 shared “Tools like Envision AI are costly and not fully accessible. Meta’s smart glasses offer good scene descriptions, but their partnership with Be My Eyes over Aira felt like privacy wasn’t a priority. I’d like to see better privacy controls before I fully rely on it.”
@@ 5.1. Scenario 1: Self Appearance: Impression Management
While the majority didn’t use GenAI for this purpose, participants shared both skepticism and cautious optimism regarding the use of GenAI tools for self-appearance management. As practice, some relied on touch and internal cues to assess their appearance, while some called friends or family for feedback as well as Be My AI or Aira to check for issues like stains on clothing. P11 illustrates how GenAI would be useful “It’s kind of uncomfortable to think about, I remember a specific time Aira saved me, my bra was showing. I was thinking of using Seeing AI.” P11 emphasized that she would consider GenAI if the tool guaranteed complete privacy, supported local face processing, and could simulate empathy or understanding, especially in tasks involving bodily presentation or vulnerability.
@@ 5.2. Scenario 2: Indoor Spatial Privacy
Participants often described managing indoor spatial privacy as important not only for personal safety but also for ensuring that their living or working spaces were socially acceptable from the perspective of others. Some participants acknowledged that GenAI could potentially streamline and augment these checks, for instance, P3 said “Now that I am really thinking about it instead of calling someone, using AI with the video would make it go faster. I would imagine GenAI integration to video streaming platforms to check clutter and particularly background checks when I am in a video call.”
@@ 5.3. Scenario 3: Sharing Visual Content in Social Media
Limitation & Expectation. Participants also discussed their expectations of GenAI which include seeking clarity on background detection, descriptive querying, identity cues, human-in-the-loop options in this scenario of sharing visual content in social media. Some also highlighted challenges in shared living spaces, where visual ambiguity arises due to overlapping personal items. As P10 explained: “I have a roommate we share space sometimes sensitive things lie around. Might not be mine, but hers. If those are in the picture, I don’t know how to manage. GenAI could help maybe.” P10 expected a way to distinguish between self vs others’ and privacy-aware tagging of visual content in a social media context. P5 expected support for identity matching so she can verify who is in the image without guessing.
@@ 5.5. Scenario 5: Visual Privacy Management by BLV Professionals
BLV Professionals expressed a tension between the accessibility benefits of GenAI tools and the risks to visual privacy, particularly when handling sensitive documents of others. Several participants working in financial, educational, and social service sectors shared workplace-specific experiences (Figure LABEL:fig2). Some described practices to protect client confidentiality. P11, a former accessibility tester “I reviewed documents with proprietary client information. We set up a closed system, we avoided bringing in third parties. Honestly, I know GenAI will be mainstream. It’s just a matter of time and the highest bidder.” Moreover, institutional professionals like P9 reported segmenting their GenAI use based on data sensitivity- “I try not to input any student documentation, grade info into Gemini right now, but I use LLMs to create material for training purposes. We are actively discussing how to institutionalize GenAI use.” A blind professional, P21, reported using generative AI tools in roles that involved handling information, such as reviewing documents to determine eligibility for benefit programs.
@@ 5.6. Scenario 6: Outdoor Spatial Privacy
In this case, many expressed use of GenAI, which are shaped by past training, current usability limitations, and broader social concerns. For instance, P13 shared skepticism toward overreliance on technology. However, she acknowledged growing community interest in tools like GoodMaps  [^5] https://goodmaps.com/, especially as their coverage of indoor spaces expands. Others echoed similar tensions. P12 described how GenAI tools such as Seeing AI felt socially awkward in public spaces: “You have to hold the phone up as the bus approaches… people can think you’re taking their photos.”
@@ 6. Discussion
6.1. Notion of Visual Privacy within GenAI.
@@
Our participants expressed mixed feelings about relying on GenAI, similar to prior research on crowd-powered accessibility tools like VizWiz (Bigham et al., 2010a). While our findings indicate the convenience of the GenAI tool in addressing long-standing challenges in managing one’s privacy independently, the impersonal nature and lack of contextual awareness of GenAI could potentially amplify privacy concerns. Participants often use GenAI for socially sensitive tasks that directly shape how others perceive them and their surroundings, or convey specific messages (e.g., competence, professionalism), even for curating the visual backdrop during video calls. These tasks carry implications for social signaling from Goffman’s theory of impression management (Tseëlon, 1992). However, the generic, decontextualized feedback from GenAI systems often failed to meet users’ nuanced expectations. For instance, our findings highlight that privacy management varies across different scenarios, while in one case, risks are centered on the user’s own visual contents, in another, risks may arise from shared environments or unintended bystanders. These distinctions underscore the importance of context and relational expectations (e.g., Nissenbaum’s contextual integrity (Nissenbaum, 2004b)).
Moreover, our findings point to systemic ambiguities in the institutional handling of GenAI use. Participants, particularly blind professionals, described being left to self-govern the use of GenAI tools for high-stakes tasks such as processing confidential disability paperwork, verifying identification, or interpreting student documents. Despite recognizing both the benefits and risks, the lack of formal guidance, secure infrastructure, or vetted tools leads users to make complex privacy decisions independently. This reflects what Ahmed et al. (Ahmed et al., 2016a) and Akter et al. (Akter et al., 2022b) have identified as structural vulnerabilities, where disabled users shoulder disproportionate responsibility in mediating privacy risks in the absence of institutional safeguards.
@@ 6.2. Design Gaps: Privacy-Preserving GenAI Ecosystems
While prior research has examined visual privacy for BLV individuals, particularly through human-mediated services like Aira and Be My Eyes (Stangl et al., 2020c; Akter et al., 2020b), our study extends this work by shedding light on how PVI users re-purpose GenAI tools for both accessibility and privacy-sensitive tasks. While existing research has explored how blind individuals use visual interpretation tools like Aira, Be My Eyes, and Seeing AI, and some emerging tools (Adnin and Das, 2024), we highlighted the design gap of GenAI tools for visual privacy management.
For example, participants preferred GenAI over human assistance for sensitive tasks like reading pregnancy tests or mammograms, yet also expressed concern over data storage (Sharma et al., 2024a) and potential model training risks (Zhou et al., 2024). These practices highlight that visual privacy involves not just what is shared, but how and with whom, introducing interpersonal and emotional dimensions of risk. In managing spatial privacy, participants emphasized complementing, rather than replacing, tactile methods with GenAI. This underscores a hybrid design space not yet central in GenAI discussions. Similarly, our findings add to research on online sharing and impression management (Park, 2020; Bennett et al., 2018) by surfacing the challenge of distinguishing between one’s own and others’ belongings in shared environment- an area yet to be integrated in mainstream GenAI tools.
@@ 6.3. Design implications
On-Device or Local Encryption. Our findings revealed a common expectation across all scenarios for local, on-device data processing, particularly for facial images, home environments, financial or medical documents, and GPS or camera inputs used by BLV individuals in both public and private space. This expectation also extended to cases where BLV users processed documents on behalf of others in professional roles (e.g., student papers, social security benefit documents for eligibility). One potential approach is to design a modular, plug-and-play GenAI model sandbox per data type: Visual, Text, Audio/Video, and Sensor Module (GPS) isolated in a trusted execution environment (e.g., Intel TDX, iOS Secure Enclave)   [^6] https://docs.trustauthority.intel.com/main/articles/concept-tees-overview.html). This technical design could be paired with audit log and explainability layers to provide privacy logs detailing accessed data, module use, and outputs. This will ensure there is no network access for the sandbox unless explicitly approved by the user.
@@ Acknowledgements.
We thank our participants for their thoughtful insights. This work was partially supported by the National Science Foundation SaTC grants (#2125925, #2148080, #2126314).

### v2

@@ 4.3. Practices of GenAI for Sensitive Content
Medical Content. Participants expressed caution when dealing with medical content while mentioning “privacy” as a reason to prefer GenAI tools over traditional practices. P16 mentioned “I avoid GenAI tools for sensitive documents like medical records. I still go for trusted humans (e.g., family or Aira agents under NDA agreements).” In contrast, P17 mentioned using Generative AI for interpreting medical records, such as mammograms and MRIs, for better explainability and privacy, “ I would call my sister and tell her to ask my brother-in-law, or make another appointment. I used Be My AI and ChatGPT to make sense of my report. Honestly, I liked the idea of being private.” Some participants used GenAI tools to read sensitive physical items. P1 shared: “I used AI for my pregnancy test. It said positive several times, and I was stressed. I took more pictures, and then it said negative. I was relieved.” This highlights how participants preferred using AI independently, even when results varied, to maintain emotional and informational privacy in personal situations.
@@
Financial Document. Participants reported adjusting their methods based on perceived sensitivity and situational privacy needs. P9 said “For my credit card, I’d rather use a bonded volunteer. but I have used Be My Eyes AI to read a handwritten check in a pinch, but I knew there was a risk—it could have exposed bank info. Sometimes I just need to know what is in the memo line [..] I don’t want to bug anyone. I try to use secure options like Aira’s free call, which is limited.” P19 on the other hand, used Seeing AI to read text on a debit card: “One time I used Seeing AI. I needed to make sure I was throwing away the old debit or new. I used it to read the card numbers and confirm which was which.” Some noted that their choice depends on the level of trust and availability.
@@
Digitized Physical Media. Participants also used GenAI tools to navigate digitized and potentially sensitive physical content. P11 expressed caution with document uploads, even on trusted platforms: “am careful with uploading documents/photos, particularly Health Equity one. Even with Aira, who has confidentiality agreements, I am picky. I’d rather call my mom or dad. I sometimes use Seeing AI as a last resort.” P20 reflected on using Microsoft Seeing AI to sort mail “Seeing AI helps me sort mail. There might be something very private in that letter, and I wouldn’t even know when I opened it with AI.” This reflects a privacy pragmatism mindset where participants recognize the potential sensitivity of what they are sharing but justify the decision based on context.
4.4. Sociopolitical Decision Making Factors in Adopting GenAI Tools
@@
We observed some decision-making factors of adopting GenAI, which is shaped by broader concerns about access, policy assumptions, and representational harms.
@@
Affordability of GenAI. Participants raised concerns about the cost of GenAI technologies and the assumptions behind their pricing. P11 echoed these concerns, criticizing the pricing and accessibility of tools like Envision AI and noting disappointment in how products are marketed and supported. P11 shared “Tools like Envision AI are costly and not fully accessible. Meta’s smart glasses offer good scene descriptions, but their partnership with Be My Eyes over Aira felt like privacy wasn’t a priority. I’d like to see better privacy controls before I fully rely on it.”
@@ 5.1. Scenario 1: Self Appearance: Impression Management
While the majority didn’t use GenAI for this purpose, participants shared both skepticism and cautious optimism regarding the use of GenAI tools for self-appearance management. As practice, some relied on touch and internal cues to assess their appearance, while some called friends or family for feedback, as well as Be My AI or Aira to check for issues like stains on clothing. P11 illustrates how GenAI would be useful “It’s kind of uncomfortable to think about, I remember a specific time Aira saved me, my bra was showing. I was thinking of using Seeing AI.” P11 emphasized that she would consider GenAI if the tool guaranteed complete privacy, supported local face processing, and could simulate empathy or understanding, especially in tasks involving bodily presentation or vulnerability.
@@ 5.2. Scenario 2: Indoor Spatial Privacy
Participants often described managing indoor spatial privacy as important not only for personal safety but also for ensuring that their living or working spaces were socially acceptable from the perspective of others. Some participants acknowledged that GenAI could potentially streamline and augment these checks, for instance, P3 said, “Now that I am really thinking about it, instead of calling someone, using AI with the video would make it go faster. I would imagine GenAI integration to video streaming platforms to check clutter, and particularly background checks when I am in a video call.”
@@ 5.3. Scenario 3: Sharing Visual Content in Social Media
Limitation & Expectation. Participants also discussed their expectations of GenAI which include seeking clarity on background detection, descriptive querying, identity cues, and human-in-the-loop options in this scenario of sharing visual content in social media. Some also highlighted challenges in shared living spaces, where visual ambiguity arises due to overlapping personal items. As P10 explained: “I have a roommate we share space sometimes sensitive things lie around. Might not be mine, but hers. If those are in the picture, I don’t know how to manage. GenAI could help maybe.” P10 expected a way to distinguish between self vs others’ and privacy-aware tagging of visual content in a social media context. P5 expected support for identity matching so she can verify who is in the image without guessing.
@@ 5.5. Scenario 5: Visual Privacy Management by BLV Professionals
BLV Professionals expressed a tension between the accessibility benefits of GenAI tools and the risks to visual privacy, particularly when handling sensitive documents of others. Several participants working in financial, educational, and social service sectors shared workplace-specific experiences. Some described practices to protect client confidentiality. P11, a former accessibility tester “I reviewed documents with proprietary client information. We set up a closed system, and we avoided bringing in third parties. Honestly, I know GenAI will be mainstream. It’s just a matter of time and the highest bidder.” Moreover, institutional professionals like P9 reported segmenting their GenAI use based on data sensitivity- “I try not to input any student documentation, grade info into Gemini right now, but I use LLMs to create material for training purposes. We are actively discussing how to institutionalize GenAI use.” A blind professional, P21, reported using generative AI tools in roles that involved handling information, such as reviewing documents to determine eligibility for benefit programs.
@@ 5.6. Scenario 6: Outdoor Spatial Privacy
In this case, many expressed use of GenAI, which is shaped by past training, current usability limitations, and broader social concerns. For instance, P13 shared skepticism toward overreliance on technology. However, she acknowledged growing community interest in tools like GoodMaps  [^5] https://goodmaps.com/, especially as their coverage of indoor spaces expands. Others echoed similar tensions. P12 described how GenAI tools such as Seeing AI felt socially awkward in public spaces: “You have to hold the phone up as the bus approaches… people can think you’re taking their photos.”
@@ 6. Discussion
6.1. Notion of Visual Privacy within GenAI
@@
Our participants expressed mixed feelings about relying on GenAI, similar to prior research on crowd-powered accessibility tools like VizWiz (Bigham et al., 2010a). While our findings indicate the convenience of the GenAI tool in addressing long-standing challenges in managing one’s privacy independently, the impersonal nature and lack of contextual awareness of GenAI could potentially amplify privacy concerns. Participants often use GenAI for socially sensitive tasks that directly shape how others perceive them and their surroundings, or convey specific messages (e.g., competence, professionalism), even for curating the visual backdrop during video calls. These tasks carry implications for social signaling from Goffman’s theory of impression management (Tseëlon, 1992). However, the generic, decontextualized feedback from GenAI systems often failed to meet users’ nuanced expectations. For instance, our findings highlight that privacy management varies across different scenarios, while in one case, risks are centered on the user’s own visual contents, in another, risks may arise from shared environments or unintended bystanders. These distinctions underscore the importance of context and relational expectations (e.g., Nissenbaum’s contextual integrity (Nissenbaum, 2004)).
Moreover, our findings point to systemic ambiguities in the institutional handling of GenAI use. Participants, particularly blind professionals, described being left to self-govern the use of GenAI tools for high-stakes tasks such as processing confidential disability paperwork, verifying identification, or interpreting student documents. Despite recognizing both the benefits and risks, the lack of formal guidance, secure infrastructure, or vetted tools leads users to make complex privacy decisions independently. This reflects what Ahmed et al. (Ahmed et al., 2016) and Akter et al. (Akter et al., 2022) have identified as structural vulnerabilities, where disabled users shoulder disproportionate responsibility in mediating privacy risks in the absence of institutional safeguards.
@@ 6.2. Design Gaps: Privacy-Preserving GenAI Ecosystems
While prior research has examined visual privacy for BLV individuals, particularly through human-mediated services like Aira and Be My Eyes (Stangl et al., 2020b; Akter et al., 2020), our study extends this work by shedding light on how BLV users re-purpose GenAI tools for both accessibility and privacy-sensitive tasks. While existing research has explored how blind individuals use visual interpretation tools like Aira, Be My Eyes, and Seeing AI, and some emerging tools (Adnin and Das, 2024), we highlighted the design gap of GenAI tools for visual privacy management.
Our findings indicated GenAI’s distinct affordances to emphasize the potential for context-aware, customizable, and conversational forms of privacy mediation, in contrast to rule-based or static vision systems. For example, participants preferred GenAI over human assistance for sensitive tasks like reading pregnancy tests or mammograms, yet also expressed concern over data storage (Sharma et al., 2024a) and potential model training risks (Zhou et al., 2024). These practices highlight that visual privacy involves not just what is shared, but how and with whom, introducing interpersonal and emotional dimensions of risk. In managing spatial privacy, participants emphasized complementing, rather than replacing, tactile methods with GenAI. This underscores a hybrid design space not yet central in GenAI discussions. Similarly, our findings add to research on online sharing and impression management (Park, 2020; Bennett et al., 2018) by surfacing the challenge of distinguishing between one’s own and others’ belongings in shared environment an area yet to be integrated in mainstream GenAI tools. Unlike rule-based tools, which operate on fixed detection pipelines, GenAI allows users to dynamically shape and refine their queries, thus embedding privacy preferences into interaction flows.
@@
6.3. Sociopolitical Barriers, Policy Implications and Risk of GenAI Censorship
Our findings highlighted that the interaction of BLV users with GenAI tools is influenced not just by technical capabilities and personal preferences, but also by sociopolitical structures. Several participants shared concerns related to the absence of explicit institutional safeguards when employing GenAI tools in situations that are seemingly in need of legal protections, such as the affordability of technology integrated with GenAI. While the Assistive Technology Act of 2004 mandates that states ensure access to assistive technologies (Bausch et al., 2005), participants shared how the high costs of advanced AI-based tools (like Envision Glasses) place an unfair burden on users.
Our findings also shed light on the restricted GenAI functionality, which may limit BLV users’ ability to access and share content. For example, over-filtering or censoring essential indicators related to appearance, gender, or race, likely influenced by platforms’ bias mitigation policies, lead to vague outputs which undermine both accessibility and utility. Under the Americans with Disabilities Act (ADA) (ADA, 1990), individuals with disabilities should have the same opportunities and access as non-disabled individuals, legislation that has been increasingly applied to digital environments, including AI tools (Neogi, 2024). However, our findings underscore that blanket policies (e.g., redacted descriptors, broad safety filters) may potentially conflict with BLV users’ needs for detailed, descriptive, and context-rich feedback.
Participants also described reliance on GenAI in understanding sensitive documents (e.g., insurance forms, student records). In professional settings, this situates BLV individuals at the intersection of ADA compliance and institutional data protection mandates such as, HIPAA, FERPA. While there are guidelines, such as, FCC’s CVAA (2010) (No, [n. d.]) for accessible communications and ADA Title I (ada, [n. d.]) for workplace equity, participants are at a crossroads without institutional policy guidance. We believe addressing these tensions needs more than guidelines; rather, fundamental rethinking of GenAI design for an enterprise-safe and privacy-preserving ecosystem that both complies with data protection while allowing BLV professionals to leverage the benefits of GenAI meaningfully.
@@ 6.4. Design implications
On-Device or Local Encryption. Our findings revealed a common expectation across all scenarios for local, on-device data processing, particularly for facial images, home environments, financial or medical documents, and GPS or camera inputs used by BLV individuals in both public and private spaces. This expectation also extended to cases where BLV users processed documents on behalf of others in professional roles (e.g., student papers, social security benefit documents for eligibility). One potential approach is to design a modular, plug-and-play GenAI model sandbox per data type: Visual, Text, Audio/Video, and Sensor Module (GPS) isolated in a trusted execution environment (e.g., Intel TDX, iOS Secure Enclave)   [^6] https://docs.trustauthority.intel.com/main/articles/concept-tees-overview.html). This technical design could be paired with audit logs and explainability layers to provide privacy logs detailing accessed data, module use, and outputs. This will ensure there is no network access for the sandbox unless explicitly approved by the user.
@@
6.5. Limitations
Our work has a number of limitations. Our participants’ sample of 21 may not be fully representative of the broader BLV community. From our study, very few participants had the affordability to use emerging assistive tools, such as Meta Rayban Glass, Envision Glass, etc, for some use case relevant to visual privacy, which might suggest limited representation of higher socio-economic groups.
@@ Acknowledgements.
We thank our participants for their thoughtful insights. This work was partially supported by the National Science Foundation SaTC grants (#2125925, #2148080, #2126314). Conflict of interest disclosure: Leah Findlater is also employed by and has a conflict of interest with Apple Inc. Any views, opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and should not be interpreted as reflecting the views, policies, or position, either expressed or implied, of Apple Inc.